OpenStudy (anonymous):

Which number has the greatest value? How do you know? 2.3 x 10-3 0.23 x 10-5 23x10-7

OpenStudy (anonymous):

2.3 x 10 - 3 = 16.1 0.23 x 10-5 = 1.15 23x10-7 = 69 (7) (5) (3)

OpenStudy (anonymous):

Ooh haha thank you